A straight aluminum wire 30m Long is subject to a tensile stress of 70MPa.
1. Determine the total elongation of the wire.
2. What temperature change would produce the same elongation
Take: Elastic modulus = 200GNM^-2 And temperature coefficient = 25 × 10^-6
1
Expert's answer
2020-06-10T07:51:15-0400

l=30 m, "\sigma_T=" 70 MPa,Y=200 GPa

we know that

stress= "Y\times strain"

"\sigma _T= Y \times \frac{\Delta l}{l}"

here , "\Delta l = elongation"

"\Delta l= \frac{70 \times 30}{200}=10.5 mm"


(b) as per the concept of thermal stress

"\Delta l= l\alpha \Delta T"

10.5"\times 10^{-3}" = "30\times 25\times 10^{-6} \Delta T"


"\Delta T =" 13.33 "^oC" This is the temperature difference which will produce above elongation